OpenStudy (anonymous):

solve for x. 1/27=9^7x+3

OpenStudy (stormfire1):

\[x=\frac{\frac{1}{27}-3}{9^7}\]

OpenStudy (anonymous):

x= (-9/14)

OpenStudy (stormfire1):

How did you get that?

OpenStudy (anonymous):

3^(-3)=3^2*(7x+3) 3^(-3)=3^(14x+6) this implies that -3=14x+6 i.e., 14x= -6-3 x=-9/14

OpenStudy (precal):

Stormfire 1 is wrong Sheg is correct

OpenStudy (stormfire1):

Still not seeing it...I see the original equation as:\[\frac{1}{27}=9^7x+3\]\[3^{-3}-3=9^7x\]\[x=\frac{3^{-3}+3}{9^7}\]Maybe I'm reading the original equation differently...

OpenStudy (precal):

The rule is you can log both sides or unlog both sides. You also, can take the powers and set them equal to each other. 9 can be written with the base of 3 just like you wrote 3 to the negative 3 power. Besides your x value does not match sheg's x value. You should be able to sub your x value back into the equation in order for it to be true. Your x value does not work.

OpenStudy (stormfire1):

I'm aware of the log & exponent rules :) As you can see from my post, I read the original equation differently. I will admit that I flipped a sign in the equation editor by mistake in my last post...which may have thrown you off. My original post did have the correct answer for the problem as I read it: \[\frac{1}{27}=9^7x+3\] However, sheg reads it as: \[\frac{1}{27}=9^{7x+3}+3\]which explains the different answers. The bottom line is that a few parentheses in the original question would have made everything more clear :)